### v2

# Aligning Large Language Models by On-Policy Self-Judgment

Abstract
Existing approaches for aligning large language models with human preferences face a trade-off that requires a separate reward model (RM) for on-policy learning. In this paper, we present a novel alignment framework, Self-Judge that (1) does on-policy learning and 2) is parameter efficient, as it does not require an additional RM for evaluating the samples for on-policy learning. To this end, we propose Judge-augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. Specifically, we view the pairwise judgment task, choosing the better response from a response pair, as a special case of the instruction-following task. The resulting model can judge preferences of on-the-fly responses from current policy initialized from itself. Experimental results show the efficacy of Self-Judge, outperforming baselines in preference benchmarks. We also show that the rejecting sampling by itself can improve performance further without an additional evaluator [^1] The code is available at github.com/oddqueue/self-judge.
[…]
1 Introduction
Research on aligning Large Language Models (LLMs) with human preference has increasingly gained attention in recent years Askell et al. (2021); Ouyang et al. (2022); Bai et al. (2022a); Rafailov et al. (2023). Reinforcement Learning from Human Feedback (RLHF) is the most dominant approach for the alignment of LLMs for human preferences Ziegler et al. (2020). It utilizes a reward model (RM) to estimate the human preference scores for the generated responses from policy. The policy is updated with on-policy Reinforcement Learning (RL) to maximize the estimated rewards of sampled responses regularized with the KL divergence between the current policy and reference policy. However, RLHF requires a complex setup for on-policy updates because of its simultaneous use of an RM with the reference model.
Figure 1: In our framework, Self-Judge, a single model is trained not only to generate responses but also to perform a judgment task, where it selects the better of the two responses through a single token prediction. This enables on-policy self-training by performing judgments on current policy for improving itself.
Figure 2: An overview of Self-Judge. 1) We train an LLM to act as a Judge Model (JM), which can both generate responses and compare response pairs. We train the JM with a SFT dataset augmented with the pairwise judgment task where the better response can be selected by a single token. 2) We initialize a policy and a fixed reference model from the trained JM. Then, the policy model samples response pairs, and the reference model performs judgments on the pairs for giving feedback with preference orders. 3) We perform a rejection sampling by a tournament on responses from the policy through the judgments by itself for further improvements at inference time.
In contrast, a line of research Rafailov et al. (2023); Azar et al. (2023) proposes discarding on-policy learning and optimizing from preference orders in static datasets without RMs (i.e., offline learning), whereas other studies propose constructing datasets with responses sampled from the policy and labeled by separated evaluators (i.e., off-policy learning) Zhao et al. (2023); Liu et al. (2023); Xu et al. (2023); Gulcehre et al. (2023); Dong et al. (2023). However, offline learning can lead to sub-optimal results due to the lack of exploration Mediratta et al. (2023), and off-policy learning could cause degeneration if an appropriate replay buffer strategy is not used Zhang and Sutton (2017).
In this paper, we present Self-Judge, a simplified on-policy training scheme that trains a single model to perform on-the-fly feedback for current policy for improving itself. To this end, we introduce Judge-augmented Supervised Fine-tuning (JSFT) to obtain a model that can both generate a response and perform pairwise comparison, as shown in Figure 1. Specifically, we regard the pairwise judgment task, choosing the better response from a response pair, as a special case of the instruction-following task, which can be answered in a single token and optionally with a rationale. Self-Judge initializes the current policy and reference policy from JSFT’s resulting model. Self-Judge samples a response pair from the current policy and chooses a better response in the pair by the reference policy, then updates the current policy with preference orders without an RM Rafailov et al. (2023); Zhao et al. (2023); Azar et al. (2023).
Experimental results show that Self-Judge outperforms RLHF and other offline and off-policy approaches Rafailov et al. (2023); Liu et al. (2023); Gulcehre et al. (2023); Dong et al. (2023) on preference benchmarks. Unlike existing approaches, Self-Judge leverages on-policy learning while not introducing an additional evaluator. The results demonstrate the effectiveness and parameter efficiency of Self-Judge performing on-policy self-training for LLM alignment. Furthermore, we show that Self-Judge can maximize performance through self-rejection that selects the best response from its own responses using its judgment capabilities learned through JSFT. In particular, the performance gains are significant when the pairwise judgment task of JSFT involves comparisons based on principles with rationale for the decision.
In summary, our main contributions are:
  • We propose a parameter-efficient on-policy learning framework, Self-Judge, with JSFT for obtaining an initial policy that can judge.
  • We analyze the efficacy of the JSFT for judgment and suggest the best practices to exploit the improved judgment ability.
  • We show resulting models from JSFT can self-improve by acting as a judge: on-policy self-training and self-rejection at inference time.
2 Preliminaries
Aligning by Reinforcement Learning
To align LLMs with human preferences, RLHF Ziegler et al. (2020) follows three stages: 1) obtaining an initial policy $\pi_{\text{ref}}$ through SFT from the pre-trained LLM, 2) training an RM from human preference triplets $(x,y_{w},y_{l})$, where $y_{w}$ is a chosen response, and $y_{l}$ is a rejected response for a given prompt $x$, 3) fine-tuning the initial policy by Proximal Policy Optimization (PPO) Schulman et al. (2017), with KL divergence between current policy $\pi$ and reference policy $\pi_{\text{ref}}$ as a regularization for the reward maximization objective. Generally, RMs are trained by Bradley-Terry model Bradley and Terry (1952), which minimizes negative log-likelihood of score difference, $\text{log}\sigma(r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l}))$, to compute a pointwise scalar reward for the response sampled from the current policy. This on-policy rollout procedure provides the frequent exploration of responses but introduces an additional training stage and memory usage due to the RMs.
Aligning from Preference Orders
It has been observed that RM is susceptible to language biases such as the response length, preferring longer responses over shorter ones Shen et al. (2023). Askell et al. (2021) suggest a language modeling loss on preferred context $(x,y_{w})$ during the training of the RM to address this issue. This approach, however, necessitates another pre-training stage of the RM. Direct Preference Optimization (DPO) Rafailov et al. (2023) and Sequence Likelihood Calibration with Human Feedback (SLiC-HF) Zhao et al. (2023) introduce objectives that can be trained by only preference orders $y_{w}\succ y_{l}$ without scalar rewards and need for RMs. However, offline learning methods optimized on static datasets inherently reach sub-optimal results compared to online learning Mediratta et al. (2023). Xu et al. (2023); Yuan et al. (2024) propose an online learning approach that iteratively constructs datasets by responses sampled from the policy. However, this off-policy approach with a large buffer size can induce performance degeneration Zhang and Sutton (2017).
3 Self-Judge
In this section, we describe our proposed alignment framework, Self-Judge, which utilizes a single model that acts as both policy and judge: sampling responses and judgment over response pairs. It can provide feedback on the response pairs sampled from the current policy for improving itself in an on-policy manner and also perform rejection sampling by itself, as depicted in Figure 2.
3.1 Judge Model
Using LLMs to evaluate responses of another LLM, LLM-as-a-judge, is shown promising Zheng et al. (2023); Bai et al. (2022b); Kim et al. (2023). Inspired by these studies, we leverage the generative pairwise evaluator, which we refer to as Judge Model (JM), for aligning LLMs with human preference Ethayarajh et al. (2023); Zhao et al. (2023); Liu et al. (2023). Unlike an RM, which produces a scalar score for a single response, JM simply chooses the better response between the two responses. Specifically, the JM $\pi$ is trained by maximizing the log-likelihood of the choice token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ corresponding ground-truth preference label for a given judgment template $\mathcal{C}$. Figure 3 illustrates an example of a judgment template.
Figure 3: An example of a judgment template $\mathcal{C}$. The judgment template asks which of the two responses is better for a given prompt and requests to select the choice token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ corresponding to the better response. Optionally, a principle for the judgment can be added to the judgment template, and the rationale can be included in the target sequence for training.
Judge-augmented Supervised Fine-tuning
Typically, the chosen response $y_{w}$ is used for target sequence of prompt $x$ in training preference dataset $\mathcal{D}$ on the SFT stage, i.e., $\mathcal{D}_{\text{SFT}}=\{(x,y_{w})|(x,y_{w},y_{l})\in\mathcal{D}\}$. In addition, we treat the pairwise judgment task as a special case of instruction-tuning, e.g., $\mathcal{D}_{\text{Judge}}=\{\big{(}\mathcal{C}(x,y_{w},y_{l}),\mathcal{J}\big{)}|(x,y_{w},y_{l})\in\mathcal{D}\}$. As a result, we train the pairwise judgment task as a response generation using the augmented dataset $\mathcal{D}^{+}_{\text{SFT}}=\mathcal{D}_{\text{SFT}}\cup\mathcal{D}_{\text{Judge}}$. We refer to the process of fine-tuning the LLMs on the augmented dataset $\mathcal{D}^{+}_{\text{SFT}}$ as Judge-augmented Supervised Fine-tuning (JSFT). With JSFT, we can obtain a JM that can not only compare pairwise preferences but also generate responses. Also, we can expect a better understanding of language-relevant features on preference comparison since the JM is trained to mimic good behavior $y_{w}$, similar to the observation found on RMs Askell et al. (2021).
Principle-aware Judgment with Rationale
One advantage of JM compared to the conventional method with RM is that the JM can adapt the judgment template $\mathcal{C}$ according to multiple aspects of human preferences, reflecting diverse principles Sun et al. (2023); Cui et al. (2023). Also, the generative nature of JM makes it easy to utilize rationale, the justification for the judgment in natural language, for more precise judgments, similar to the observations on instruction-tuning of LLMs Mukherjee et al. (2023). For a given principle $p\in\mathcal{P}$ where the $\mathcal{P}$ is a principle set, we devise different judgment templates $\mathcal{C}_{p}$ for each $p$. Also, we can optionally include a rationale $\mathcal{R}$ in the target sequence for the judgment task when it is available. The JM first produces a choice token $\mathcal{J}\in\{\mathcal{A},\mathcal{B}\}$ which corresponds to the ground-truth label of the better response, then rationale $\mathcal{R}$ is followed. Formally, $\pi(\mathcal{J},\mathcal{R}\mid\mathcal{C}_{p}(x,y_{w},y_{l}))$. As a result, principle-aware judgment can be conducted by simply adjusting the judgment template.
Judging Self-Generated Responses
Since JM is trained by JSFT, it can judge its own response by playing the roles of the policy and judge: this is the intuition behind the name Self-Judge. As a policy, the JM first samples responses $y_{a},y_{b}$ for a given $x$. Then, the JM, as a judge, chooses a better response between the two responses, $\pi(\mathcal{J}\mid\mathcal{C}(x,y_{a},y_{b}))$. More precisely, it averages the likelihoods of corresponding choice tokens, $\mathcal{J}$, utilizing a position-swapped judgment template to mitigate position bias Chiang et al. (2023). From the relative likelihoods of these choice tokens for each response, a pseudo-preference triplet label $(x,\hat{y}_{w},\hat{y}_{l})$ for the self-training and self-rejection can be deduced. For the principle-aware judgments, we can check the winning rate across principles or the mean likelihood of choice tokens across principles when a tie happens.
3.2 Self-Training by On-Policy Judgment
Ethayarajh et al. (2023) uses JM’s predicted likelihood of label tokens as the reward for on-policy evaluation in RLHF framework, comparing with a blank baseline response 222huggingface.co/stanfordnlp/SteamSHP-flan-t5-xl. This approach has a significant limitation; the likelihood inferred by a language model cannot be regarded as confidence without calibration Zhou et al. (2023); Zhu et al. (2023). Conversely, to leverage JMs properly, we adopt objectives that perform optimizations on preference orders since these objectives do not require pointwise scalar scores Rafailov et al. (2023); Zhao et al. (2023); Azar et al. (2023). In our framework, we regard the reference policy $\pi_{\text{ref}}$ as a judge to evaluate the preference order between generated samples from current policy $\pi_{\theta}$, as follows:
$\begin{split}&y_{a}\sim\pi_{\theta}(\cdot\mid x),y_{b}\sim\pi_{\theta}(\cdot\mid x),\\ &(x,\hat{y}_{w},\hat{y}_{l})\leftarrow\pi_{\text{ref}}(\mathcal{J}\mid\mathcal{C}(x,y_{a},y_{b})).\\ \end{split}$
That is, a single JM with JSFT is initialized for both $\pi_{\theta}$ and $\pi_{\text{ref}}$, but the latter one is frozen for the likelihood normalization Liu et al. (2023) and on-policy judgments. This setup enjoys on-policy learning without introducing an additional model for the evaluation as RLHF. If we adopt the DPO objective, the following loss is used for optimization, where $\sigma$ is the sigmoid function, and $\beta$ is a coefficient for KL divergence regularization,
$\begin{split}-\mathbb{E}\big{[}\log\sigma\big{(}\beta\log\cfrac{\pi_{\theta}(\hat{y}_{w}|x)}{\pi_{\text{ref}}(\hat{y}_{w}|x)}-\beta\log\cfrac{\pi_{\theta}(\hat{y}_{l}|x)}{\pi_{\text{ref}}(\hat{y}_{l}|x)}\big{)}\big{]}.\\ \end{split}$
3.3 Self-Rejection by Tournament
RMs can also be utilized at inference time for rejection sampling such as Best-of-$N$ sampling Stiennon et al. (2020), selecting the best response according to the reward score among oversampled $N$ responses. However, this approach requires a separate RM in addition to the policy model during inference. In contrast, JM trained within Self-Judge does not require an additional RM for evaluating responses, as it can evaluate the self-generated responses by itself.
However, judging all the possible comparison pairs to get average winning rates requires $\mathcal{O}(N^{2})$ forward passes for (${N\choose 2}$ comparisons. Thus, we adopt the tournament Zhao et al. (2023); Liu et al. (2023) for rejection sampling at inference time. We construct a tournament tree whose leaf nodes are sampled responses, and the non-leaf nodes are chosen by the winner on judgment between the child nodes. Since the tournament tree has less than $N-1$ non-leaf nodes, we can find the best response by $\mathcal{O}(N)$ forward passes, the same as the Best-of-$N$ sampling with a separated RM.
4 Experimental Setup
Method Policy Evaluator On-Policy Learning On-Memory Parameters AlpacaEval (Win) VicunaEval (Win) MT-Bench (Score) SFT $p$ 24.75 50.00 4.63 DPO SFT $2p$ 35.14 60.63 4.73 RSO SFT JM $2p$ 34.27 64.38 4.42 ReST SFT RM $p$ 27.43 55.00 4.53 RAFT SFT RM $p$ 32.50 59.38 4.43 RLHF SFT RM $3p$ 33.46 53.75 4.29 Self-Judge (Ours) JM $2p$ 44.88 76.25 4.80
Table 1: Evaluation results of models trained on HH-Helpful Bai et al. (2022a). The best result and second best result on each benchmark are represented as bold and underline. We report theoretical memory usage of model parameters required for each method where $p$ denotes the number of parameters of the base model. We use base and online splits but used the online split only for constructing the training instances of SFT. Self-Judge outperforms baselines on all benchmarks with a single JM, which can act as both policy and judge.
In this section, we discuss the experimental setup for validating our proposed framework, Self-Judge. We use two datasets: Anthropic-HH Yuan et al. (2023) and UltraFeedback Cui et al. (2023). In the experiments, we choose DPO Rafailov et al. (2023), RSO Liu et al. (2023), ReST Gulcehre et al. (2023), RAFT Dong et al. (2023), and RLHF Ziegler et al. (2020) as baselines for comparing with Self-Judge. We evaluate the resulting models from each experiment by AlpacaEval Li et al. (2023), VicunaEval Chiang et al. (2023), and MT-Bench Zheng et al. (2023). The implementation details are in Appendix A.
4.1 Datasets
Anthropic-HH
Anthropic-HH Bai et al. (2022a) is a human preference dataset on 170k dialogues, which consists of two subsets, HH-Helpful and HH-Harmless, which are labeled by the helpfulness and harmlessness principle. We focus on the HH-Helpful to better isolate and understand the benefits of Self-Judge due to the conflicting nature of helpfulness and harmlessness principles Bai et al. (2022a). HH-Helpful contains various data splits corresponding to the development stages of an AI assistant. We use the base split and include the pair $(x,y_{w})$ from the online split in the SFT dataset for analyzing transition effects on JSFT.
UltraFeedback
As obtaining human-labeled feedback is costly, utilizing AI feedback is a widely investigated alternative Bai et al. (2022b); Sun et al. (2023); OpenAI (2023). UltraFeedback Cui et al. (2023) is one of the datasets that consists of AI feedback where GPT-4 OpenAI (2023) rates responses obtained from four different language models for the 64k prompts, based on four principles of helpfulness, instruction-following, truthfulness, and honesty. Each rating contains the rationale obtained from GPT-4, which represents an explanation for the quality and rating of the corresponding response based on the given principle. We randomly split 10% of the prompts of the dataset and use them as a test set for further analysis.
4.2 Baselines
We choose DPO Rafailov et al. (2023) as an offline learning baseline and use DPO objective for self-training in Self-Judge. We include RSO Liu et al. (2023) with the DPO objective as an off-policy learning baseline, which utilizes the JM as a separate evaluator. We include baselines which utilize RMs, ReST Gulcehre et al. (2023) and RAFT Dong et al. (2023) for off-policy approach and RLHF Ziegler et al. (2020) for on-policy approach. We choose LLaMa-2-7B Touvron et al. (2023) as a base model for all experiments.
4.3 Evaluations
We evaluate the resulting models based on the three benchmarks, AlpacaEval Li et al. (2023), VicunaEval Chiang et al. (2023), and MT-Bench Zheng et al. (2023). AlpacaEval is an alignment benchmark that compares the quality of two responses on 805 questions sampled from a diverse dataset, rated by GPT-4 OpenAI (2023). We use the text-davinci-003 Ouyang et al. (2022) as the baseline model for measuring winning rates. VicunaEval is another benchmark utilizing GPT-4 as a judge for comparing two models’ responses from 80 questions on various topics. We use the SFT model on each dataset as the baseline model for measuring winning rates. MT-Bench is a multi-turn benchmark consisting of 80 instances of 2-turn questions from 8 different domains, which is evaluated by GPT-4 on a scale of 1 from 10. We report the average scores obtained on each turn.
5 Experimental Results
5.1 Main Results
Self-Judge is a strong alignment method.
In Table 1, we can see that Self-Judge outperforms all baselines. Notably, Self-Judge consistently shows the highest performance on all three benchmarks while not introducing additional parameters compared to DPO, which is trained in an offline setting. In addition, Self-Judge also shows significant strength compared to all of the baselines utilizing separate evaluators for off-policy and on-policy learning Liu et al. (2023); Gulcehre et al. (2023); Dong et al. (2023); Ziegler et al. (2020). Qualitative examples can be found in Appendix B.
5.2 Analysis of Judge Models
We conduct an ablation study with HH-Helpful to verify two hypotheses about Self-Judge: 1) JSFT improves JM’s judgment ability, and 2) JM is more compatible with direct preference optimization than RL. To this end, we train three JMs with different strategies: 1) training solely on judgment task induced from base split, 2) JSFT on base split, and 3) JSFT on base split with additional SFT data from online split. We report the performance of each JM as a policy and a judge. We further examine the resulting models on RLHF regarding the predicted likelihood of label tokens on JM as rewards similar to Ethayarajh et al. (2023). We also investigate how different sampling strategies in self-training influence the final performance.
JSFT improves JM’s judgment ability.
Table 2 shows the results regarding the first hypothesis. We can see that JM has only a marginal difference in prediction accuracy on the test split of HH-Helpful compared to RM when it is not trained by JSFT. From this observation, we confirm that the transition effects of imitation learning to judgment task occur when the judgment task is trained with canonical SFT tasks. On the other hand, including the online split for JSFT significantly improves the performance as a policy, but performance as a judge is slightly decreased compared to the JM that is only trained on the base split. We conjecture that the distribution gap between online split and base split influences the transition effects, as online split is obtained from a language model already aligned with human preferences.
Type JSFT (+ $\mathcal{D}_{\text{SFT}}$) Judge (Accuracy) Policy (Win) RM 66.11 JM 66.49 + base 68.32 5.78 + base / online 67.84 20.26
Table 2: Prediction accuracy on test split of HH-Helpful and winning rates on AlpacaEval using JM as a judge or a policy. JSFT improves not only the performance as a policy but also the performance as a judge of JM.
On-policy learning with preference orders is the best strategy for JM.
In Footnote 4, we observe that RLHF with JM as an evaluator results in a lower performance compared to conventional RM producing scalar rewards, which implies that the choice token likelihood obtained from a judgment could not be regarded as a pointwise preference score on the Bradley-Terry model Bradley and Terry (1952). Table 4 shows that on-policy learning outperforms offline and off-policy approaches using a single JM on self-training. Therefore, we can conclude that on-policy learning with preference orders, such as DPO objective, is the best strategy for JM in order to leverage JM’s superior performance on judgments compared to the RM.
Policy Evaluator JSFT (+ $\mathcal{D}_{\text{SFT}}$) AlpacaEval (Win) SFT RM 33.46 JM 26.18 + base 29.63 + base / online 28.46
Table 3: Prediction accuracy with different evaluator types. With JMs, we regard the likelihood of choice token comparing with chosen response $y_{w}$ as a reward [^4] We found that using the blank response fails on the assessment of response since the reward is always close to 1.. JM’s token likelihoods are not appropriate for the pointwise reward function in RLHF.
Type AlpacaEval (Win) VicunaEval (Win) MT-Bench (Score) Offline 28.57 58.75 4.68 Off-Policy 32.03 64.38 4.59 On-Policy 44.88 76.25 4.80
Table 4: Effect of learning strategy on self-training. On-policy learning yields the best performance.
5.3 Effects of Principle and Rationale
In this subsection, we compare JMs trained by three different JSFT strategies to verify the effect of principle-aware judgment and rationale through UltraFeedback Cui et al. (2023). The three different strategies include 1) judgment derived from overall scores across the principles, 2) principle-aware judgments, and 3) principle-aware judgments with rationales. We check the performance of each JM as a policy and a judge. We also examine how principle-aware judgment and rationale affect the performance as a judge after a self-training stage. We further report the statistics of selected responses on self-rejection according to the sampled number of responses. Additionally, we investigate the feasibility of iterative training, which regards the JM after the self-training as the initial policy and conducts several iterations of self-training.
Type P R Judge (Accuracy) Policy (Win) RM 79.9 JM 80.5 67.4 JM-P 81.6 59.3 JM-PR 84.0 65.7
Table 5: Performance as a judge or a policy by test split of UltraFeedback and AlpacaEval according to the usage of principle (P) and rationale (R) for pairwise judgment task. Principle-aware judgment with rationale (JM-PR) boosts the performance as a judge while slightly sacrificing the ability as a policy.
Figure 4: Winning rates, average response lengths, and 4-gram repetitions on AlpacaEval according to the number of sampling ($N$) for self-rejection on JM and JM-PR after self-training. Even though LLM-as-a-judge tends to favor verbose responses Zheng et al. (2023), JM-PR reliably improves winning rates as $N$ increases, with smoother increments of response lengths and lower repetitions compared to JM.
Involving principles and rationale improves JM’s performance as a judge but not as a policy.
Table 5 shows that JM’s performance as a judge increases when the JM is trained for principle-aware judgment (JM-P), and it increases more when the rationale is also used for training (JM-PR). This confirms that the pairwise judgment task can be treated as instruction, inheriting the benefits of instruction-tuning. However, we observe that degeneration in JM’s performance as a policy occurs when JM is trained solely on principle-aware judgment but recovered when JM is trained with rationale. We presume this trade-off comes from the bias in task distribution caused by increased pairwise judgment task instances for training principle-aware judgment. However, we speculate that rationale can mitigate performance degradation caused by response distribution mismatch between the pairwise judgment task and other tasks.
JM-PR excels in self-improvement.
In Table 6, we can observe that JM-PR achieves comparable performance on benchmarks even though JM-PR shows the degradation as an initial policy model in Table 5. When self-rejection is applied for JM-PR, the winning rate on AlpacaEval is reliably improved up to 8.4% with shorter response lengths and fewer repetitions compared to JM as Figure 4 although LLM-as-a-judge is likely to prefer longer responses Zheng et al. (2023). We further experiment with an iterative training scheme that regards the JM-PR after self-training as an initial policy. We observe that the performance as a policy improves while the capability as a judge is maintained, as shown in Figure 5. We find that the increase in performance as a policy diminishes with more iterations, but the performance degeneration as an initial policy can be overcome with iterative training.
Method AlpacaEval (Win) MT-Bench (Score) SFT 68.99 5.08 DPO 80.25 5.74 RSO 77.67 5.69 ReST 72.95 5.47 RAFT 71.08 5.32 RLHF 71.68 5.40 Self-Judge (JM) 80.75 6.00 + Self-Rejection ($N$ = 16) 84.47 6.08 Self-Judge (JM-PR) 79.98 6.12 + Self-Rejection ($N$ = 16) 88.39 6.14
Table 6: Evaluation results of models trained on UltraFeedback by AlpacaEval and MT-Bench. The best result and second best result without a rejection sampling are represented as bold and underline. Note that applying self-rejection on resulted models from Self-Judge does not require a separate evaluator.
Figure 5: Result of iterative self-training on AlpacaEval using JM-PR. Performance as a policy increases as iterations proceed without losing the capacity as a judge for applying self-rejection.
6 Related Work
Learning from Preference Scores
There are several approaches utilizing an RM for alignment learning. RLHF Ziegler et al. (2020) utilizes an RM for on-policy reinforcement learning. RRHF Yuan et al. (2023) maximizes the margin of log-likelihood by the rank of responses determined by the score from RM and human annotators. RAFT Dong et al. (2023) and ReST Gulcehre et al. (2023) apply rejection sampling on sampled responses through the RM to perform self-imitation learning. SALMON Sun et al. (2023) trains LLMs to generate scores for responses through principle-driven synthetic preference data utilizing the SFT model. However, all these approaches require a separate RM for the alignment procedure.
Optimizing on Preference Orders
From preference orders in the static dataset, DPO Rafailov et al. (2023) optimizes LLMs by implicit rewards without a separated RM. IPO Azar et al. (2023) proposes a modified objective using an unbounded preference mapping function to mitigate overfitting on deterministic preferences in the dataset. PCO Xu et al. (2023) utilizes cringe loss for optimization, which considers the token-level likelihood of rejected samples as contrastive training. SPIN Chen et al. (2024) performs iterative training considering the distribution gap between SFT datasets and the model’s responses as preference orders. Self-Rewarding Language Models Yuan et al. (2024) train LLMs to generate scores for a given response by chain-of-thought reasoning to construct preference datasets by self-generated responses. All these approaches differ from our work in that they do not perform on-policy learning.
Generative Pairwise Evaluator
The generative pairwise evaluator, which we refer to as JM, has been utilized in previous approaches to alignment learning. ILF Scheurer et al. (2023) selects the response that reflects human-requested feedback through JM. SLiC-HF Zhao et al. (2023) constructs a static preference dataset with responses obtained from the SFT model ordered by JM. RSO Liu et al. (2023) approximates the optimal policy of the RLHF objective with rejection sampling through JM’s choice token likelihood. OAIF Guo et al. (2024) adopts pre-aligned LLMs to perform the pairwise judgment task for on-policy learning, not by fine-tuning to JM. All of these approaches focus on utilizing a separate evaluator and do not address self-improvement.
7 Discussion
Recent works such as Self-Rewarding Language Models Yuan et al. (2024) and OAIF Guo et al. (2024), which are optimized with pairwise preferences, have intersections with Self-Judge in aspects of self-training and on-policy sampling. However, Self-Rewarding Language Models could not provide on-policy feedback because evaluations are performed through chain-of-thought reasoning. Also, OAIF does not address self-training since it employs a pre-aligned language model. Table 7 illustrates the difference between these approaches compared to Self-Judge.
Method Self-Training On-Policy Self-Rewarding LMs Yuan et al. (2024) OAIF Guo et al. (2024) Self-Judge (Ours)
Table 7: Comparison between Self-Judge and recent concurrent works on preference alignment learning.
8 Conclusion
We propose a parameter-efficient on-policy preference alignment framework, Self-Judge, introducing Judge-augmented Supervised Fine-tuning (JSFT). A model trained by JSFT can perform feedback to the current policy for improving itself by acting as a judge. This self-training does not require additional training stages and parameters for a reward model during the policy updates. Our resulting model outperforms RLHF, offline, and off-policy baselines in preference benchmarks, demonstrating the advantages of on-policy learning and parameter efficiency of Self-Judge. Besides, we provide various analyses on the best configurations and efficacy of the proposed JSFT. Specifically, JSFT boosts performance as a judge, and involving comparisons based on principle with rationale about decision leads to further improvement. This enhanced judging capability leads to further self-improvement as a policy at inference time by self-rejection over the model’s own responses.
Limitations
To achieve parameter-efficient on-policy self-training, Self-Judge assumes the presence of a human preference dataset of examples for pairwise judgment tasks on JSFT. Therefore, if human preference datasets Bai et al. (2022a) or strong teacher models for constructing AI Feedback datasets Cui et al. (2023) are not available, Self-Judge can not be utilized. This means that Self-Judge has a limitation compared to self-alignment approaches that can construct a preference dataset when there is no preference dataset at all Bai et al. (2022b); Sun et al. (2023); Yuan et al. (2024). Additionally, our experiments do not focus on safety. Thus, using Self-Judge without reviewing safety guards may lead to potentially inappropriate responses.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) Grant funded by the Korea government(MSIT), Artificial Intelligence Graduate School Program, Yonsei University, under Grant 2020-0-01361. The authors thank the members of NAVER Cloud and Yonsei University for their constructive comments. We are also grateful to Seungju Han for valuable discussions and helpful feedback on earlier drafts of this paper.
Appendix A Implementation Details
A.1 Pre-processing of Datasets
Target HH-Helpful UltraFeedback SFT 65842 57569 Preference 43835 57266 (+ Principle) 202887
Table 8: The number of training examples in each dataset according to the target datasets.
For the HH-Helpful555huggingface.co/datasets/Anthropic/hh-rlhf, MIT License, Copyright (c) 2022 Anthropic Bai et al. (2022a), we parse only the content of each turn through the role header in the dataset itself. During this process, if a role header exists redundantly (e.g., Human: Assistant: [content]), we remove all subsequent headers. We perform the roll-out procedures from the last assistant turn in each dialogue. For the UltraFeedback666huggingface.co/datasets/openbmb/UltraFeedback, MIT License, Copyright (c) 2023 THUNLP Cui et al. (2023), we use the mean score across principles as the overall rank of the responses. We choose the longer response as a better response when we have a tie Hosking et al. (2023). We randomly sample one response as a rejected response $y_{l}$ that is inferior in rank or score on each principle for principle-aware judgment. When a rationale is utilized in training, we remove responses that include a comparative explanation against other responses. Table 8 shows the number of resulted training examples on each dataset after the described pre-processing.
A.2 Hyperparameters
Hyperparameters Initial Feedback Epoch 1 3 Batch Size 128 64 Learning Rate 2e-5 5e-6 LR Scheduler cosine constant Warm-up Ratio 0.03 0.1 Temperature 1.0 Top-p 0.9 Max New Tokens 768 Optimizer AdamW $(\beta_{1},\beta_{2})$ (0.9, 0.999) Gradient Clipping 1.0 Max Sequence Length 2048           LoRA (r, $\alpha$) (8, 16) Dropout 0.1 RLHF Mini-batch Size 32 Inner Epochs 1 KL Scheduler (0.2, 6.0) ReST $\tau$ [0.7, 0.8, 0.9] DPO, RSO, Self-Judge $\beta$ 0.1
Table 9: Hyperparameters of the experiments. Initial refers to the value of hyperparameters for SFT, RM, and JM. Feedback refers the value of hyperparameters for baselines and Self-Judge.
We apply ReST Gulcehre et al. (2023) with $G=1,I=3$, regarding a single step of Improve step as one epoch of training and $\tau$ as a quantile threshold on reward distribution. We sample 8 responses per prompt for RAFT Dong et al. (2023) and RSO Liu et al. (2023). We randomly choose one of the sampled responses as a baseline response for JMs and accepted a maximum of 1 response per prompt for RSO. We do not conduct a hyperparameter search. Table 9 shows hyperparameters used in the experiments.
A.3 Training Details
We perform full fine-tuning to obtain the initial policy and evaluators, RM or JM initializing from LLaMA-2-7B777huggingface.co/meta-llama/Llama-2-7b, LLAMA 2 Community License, Copyright (c) 2023 Meta Platforms Touvron et al. (2023). We apply LoRA Hu et al. (2021) for computational efficiency of fine-tuning the baselines and Self-Judge. We calculate language modeling loss on responses of assistant for SFT and sequences after the judgment template $\mathcal{C}$ for the judgment task. We train JM with both position-swapped comparisons on the judgment template for given triplet $(x,y_{w},y_{l})$ to avoid position bias Zheng et al. (2023). We use the prompt $x$ used on the pairwise preference comparison for roll-out procedures. Every experiment is conducted with a maximum of 8 NVIDIA A100 80GB VRAM GPUs for up to 90 hours. The baselines are implemented based on the Transformers888github.com/huggingface/transformers, PEFT999github.com/huggingface/peft and TRL101010github.com/huggingface/trl library.
A.4 Evaluation Details
We use the checkpoints from the end of training for evaluation without any model selections on every method. Since VicunaEval Chiang et al. (2023) and MT-Bench Zheng et al. (2023) contain question types that are forced to use greedy decoding in the pre-defined configuration for each benchmark, we only apply the self-rejection on questions that do not require the greedy decoding on evaluation.
A.5 Prompt Templates
[…]
Table 10: The dialogue template for the experiments.
HH-Helpful Below are a series of dialogues between a human user and an AI assistant. The AI tries to be helpful, harmless, and honest. UltraFeedback Below are a series of dialogues between a human user and an AI assistant. The AI tries to be helpful, honest, truthful and follows the instructions from human.
Table 11: The default system messages for each dataset.
Table 10 shows the dialogue template used for the experiments and Table 11 shows default system messages used for each dataset. We use the default system message for both training and inference on HH-Helpful. In the case of UltraFeedback, we use a custom system prompt for the chosen response in the dataset when constructing training examples. We use the default system prompt for training on the judgment task and inference. For the judgment template $\mathcal{C}$, we use the template shown in Table 12 for HH-Helpful and the template shown in Table 13 for UltraFeedback. For the principle-aware judgments on UltraFeedback, we modify the guidance about principles as a system prompt as shown in Table 14  Cui et al. (2023). We use the principle-aware judgment template $\mathcal{C}_{p}$ shown in Table 15 and slightly modify this when rationale $\mathcal{R}$ is included in the training, as shown in Table 16.
[…]
Table 12: The judgment template $\mathcal{C}$ for HH-Helpful. The target sequence for the training judgment task is in bold. The default system message is shown in Table 11.
<|system|> {default system message} <|user|> Which is the better response to be an assistant who is helpful, honest, truthful and following the given instruction from user? Instruction: ------- {prompt} ------- Response A: ------- {response A} ------- Response B: ------- {response B} ------- Please choose either A or B. <|assistant|> Sure! The option which is more helpful, honest, truthful and following the given instruction from user would be Response {A or B}
Table 13: The judgment template $\mathcal{C}$ for UltraFeedback. The target sequence for the training judgment task is in bold. The default system message is shown in Table 11.
Helpfulness Under the principle of ’helpfulness’, the assistant should provide users with accurate, relevant, and up-to-date information, ensuring that the content is positive, interesting, engaging, educational, and helpful. Honesty Under the principle of ’honesty’, the assistant should be honest about whether it knows the answer and express its uncertainty explicitly. The assistant should confident on questions it knows well and be modest on those it is unfamiliar with using weakeners such as ’I guess’, ’I suppose’, ’probably’, and ’perhaps’ to express uncertainty. Instruction Following Under the principle of ’instruction following’, the assistant should align the output with intent of instruction, by understanding the task goal (intended outcome) and restrictions (text styles, format or designated methods, etc.). Truthfulness Under the principle of ’truthfulness’, the assistant should answer truthfully and be faithful to factual knowledge as well as given contexts, never making up any new facts that aren’t true or cannot be grounded in the instruction.
Table 14: The principle-aware system messages for UltraFeedback.
<|system|> {principle-aware system message} <|user|> Which is the better response for an assistant when following the principle of ’{principle}’ for a given instruction? Instruction: ------- {prompt} ------- Response A: ------- {response A} ------- Response B: ------- {response B} ------- Please choose either A or B according to the principle of ’{principle}’. <|assistant|> Sure! The option which is better guided by the principle of ’{principle}’ would be Response {A or B}
Table 15: The principle-aware judgment template $\mathcal{C}_{p}$ for UltraFeedback. The target sequence for the training judgment task is in bold. The principle-aware system messages are shown in Table 14.
<|system|> {principle-aware system message} <|user|> Which is the better response for an assistant when following the principle of ’{principle}’ for a given instruction? Instruction: ------- {prompt} ------- Response A: ------- {response A} ------- Response B: ------- {response B} ------- Please choose either A or B and evaluate each response according to the principle of ’{principle}’. <|assistant|> Sure! Here is the decision of better response and evaluations of each response, guided by the principle of ’{principle}’. Decision: Response {A or B} Evaluation of Response A: {rationale A} Evaluation of Response B: {rationale B}
Table 16: The principle-aware judgment template $\mathcal{C}_{p}$ when rationale $\mathcal{R}$ is included in UltraFeedback. The target sequence for the training judgment task is in bold.
Appendix B Qualitative Examples
We compare responses of the resulting models from baseline and Self-Judge trained on HH-Helpful. We further compare the difference after self-training between JM and JM-PR on UltraFeedback by the results of self-rejection to understand the effects of principle-aware judgment and rationale. In Table 17, we observe that Self-Judge provides richer information requested by the prompt compared to other baselines. Furthermore, we confirm that JM is likely to select a response that contains unnecessarily verbose sentences as the best response compared to JM-PR, as shown in Table 18.
Prompt what is the name of chris tucker first movie
DPO The name of Chris Tucker’s first movie is “Friday”. It was released in 1995. The movie was written by Tucker and Ice Cube, and also stars Ice Cube and Nia Long. RSO Chris Tucker’s first major film role was in the 1995 comedy film “Friday”. He also had roles in “The Fifth Element” and “Money Train”. His most famous films include “Rush Hour”, “The Best Man”, and “Rush Hour 2”. ReST The name of Chris Tucker’s first movie was “Friday”. RAFT The name of Chris Tucker’s first movie is “Friday”, which was released in 1995. RLHF His first film was "Friday", released in 1995. The film was a great success, and helped launch him to fame. It earned him a Golden Globe nomination for Best Actor in a Musical or Comedy, and received critical acclaim. It is considered a classic modern comedy movie, and has become a cult film. A sequel was released in 2007, and a third film in the franchise is planned for 2017. Self-Judge The first movie that Chris Tucker starred in was the 1995 comedy film “Friday”, directed by F. Gary Gray. The film is a classic urban comedy, centering around the adventures of two friends, played by Tucker and Ice Cube, as they spend the day hanging out in their Los Angeles neighborhood. The film was a breakout hit for Tucker, establishing him as a prominent comedic actor, and was followed by a successful sequel in 1997. Tucker then went on to star in several major Hollywood films, including the hit comedy films “Boomerang” and “Tommy” as well as the drama “Rush Hour” and its sequels. He has also had a successful career as a stand-up comedian, performing sold-out shows worldwide.
Table 17: The response examples of resulting models from each method on AlpacaEval.
Prompt Take a look at the contract and extract the parties of the agreement from it. As of the 13th day of December 2021, this Music Recording Agreement ("Agreement") is made between Good Kid, a Toronto-based musical group ("Artist"), and Universal Music Group, a record label with license number 545345 ("Record Label"). The Artist and the Recording Label may be referred to in this Agreement individually as "Parties" and collectively as "Parties." Work under this Agreement will begin on March 15, 2022, and will last for the duration of the Agreement.
JM ($\mathbf{N}$ = 1) The parties of the Music Recording Agreement are: 1. Good Kid, a Toronto-based musical group ("Artist") 2. Universal Music Group, a record label with license number 545345 ("Record Label") Both parties are referred to individually as "Parties" in the Agreement and collectively as "Parties." This means that the Artist and the Record Label are the two main participants in the agreement. JM-PR ($\mathbf{N}$ = 1) The parties involved in the Music Recording Agreement are: * Good Kid, a Toronto-based musical group, represented by the artist. * Universal Music Group, a record label with license number 545345, represented by the record label.) This agreement is effective as of the 13th day of December 2021 and will last until the expiration of the agreement. The work under this agreement will begin on March 15, 2022, and will continue for the duration of the agreement.
JM ($\mathbf{N}$ = 16) Based on the provided Music Recording Agreement, the parties involved in the agreement are: 1. Good Kid, a Toronto-based musical group (Artist). 2. Universal Music Group, a record label with license number 545345 (Record Label). The Agreement refers to the Artist and the Record Label as "Parties" and collectively as "Parties" throughout the document. Therefore, the parties to the Music Recording Agreement are: * Good Kid, a Toronto-based musical group (Artist) * Universal Music Group, a record label with license number 545345 (Record Label) The Agreement outlines the terms and conditions of the recording agreement between the Artist and the Record Label, including the commencement date, the duration of the agreement, and the scope of work. JM-PR ($\mathbf{N}$ = 16) Based on the provided contract, the parties involved in the Music Recording Agreement are: 1. Good Kid, a Toronto-based musical group ("Artist") 2. Universal Music Group, a record label with license number 545345 ("Record Label") These two parties are involved in the agreement and are referred to individually as "Parties" and collectively as "Parties" throughout the document.
Table 18: The response examples of JM and JM-PR by applying self-rejection on AlpacaEval.